A bag contains​ pennies, nickels,​ dimes, and quarters. There are 50 coins in all. Of the​ coins, 10​% are pennies and 40​% are dimes. There are 9 more nickels than pennies. How much money does the bag​ contain?



Answer by ikleyn(52788)

.
10% of 50 coins is/are 5 coins.  So, 5 coins are pennies.


40% of 50 coins is/are 20 coins.  So, 20 coins are dimes.


The number of nickels is  (5 + 9) = 14, according to the condition.


Then quarters are the rest (50 - 5 - 20 - 14) = 11 coins.


Now you can easily calculate the total and answer the question


    Total = 5*1 + 14*5 + 20*10 + 11*25 = 550 cents = $5.50.     ANSWER

Solved.